what times 4/3 equals 1/3
Answer by unlockmath(1688)

Hello,
This (what times 4/3 equals 1/3) can be rewritten as:
4/3(x)=1/3
Multiply by 3 to get:
4x=1
Divide by 4 and we have:
x=1/4
Make sense?
RJ
www.math-unlock.com